OpenStudy (anonymous):

Find the slope of the line passing through the points and . (2,-4) and (7,5)

OpenStudy (anonymous):

When finding the slope of a line (m) passing through 2 points (-9, -6) and (-9, -8) we can use the slope formula. \[m = \frac{ Y2- Y1 }{ X2-X1 }\] Lets use point (2,-4) as (x1,y1) and point(7,5) as (x2,y2). \[m = \frac{ 5 - (-4) }{ 7 - 2 }\] which further becomes \[m = \frac{ 9 }{ 5 }\] which is the slope